OpenStudy (anonymous):

At an accident scene on a level road, investigators measure a car's skid marks to be 84.9m long. The accident occurred on a road with a coefficient of friction estimated to be .34. What was the speed of the car as the driver slammed and locked the brakes?

OpenStudy (anonymous):

1/2 x m x v^2 = Displacement x m x g(gravity) x u(coefficient of friction) > the masses cancel out > .5v^2 = Dgu >plug n' chug: v^2 = (2)(84.9m)(p.81m/s^2)(.34) >v^2 = 565.77 v = 565.77^.5 v = 23.79m/s I figured it out. Thought i might as well share!

OpenStudy (amistre64):

\[F_k=\mu_k~\vec n~:~F=ma\] \[F_k=\mu_k~mg~:~F=ma\] \[\mu_k~mg=ma\] \[\mu_k~g=a\] \[v_f^2=v_o^2+2a\Delta x\] \[v_f^2=v_o^2+2\mu_k~g~\Delta x\] seems about fair to me

OpenStudy (amistre64):

solve for Vo

OpenStudy (amistre64):

and since acceleration is slowing , the negative aint an issue \[v_f^2=v_o^2+2\mu_k~g~\Delta x\] \[v_f^2-2\mu_k~g~\Delta x=v_o^2\] \[\sqrt{v_f^2-2\mu_k~g~\Delta x}=v_o\] \[\sqrt{2(.34)(9.8)(84.9)}=v_o=23.79\]

OpenStudy (amistre64):

yeah, we get to the same results

OpenStudy (anonymous):

YAY!